OpenStudy (anonymous):

True or false: The quantity represented by theta_0 is a function of time (i.e., is not constant).

OpenStudy (anonymous):

Theta_0 ? is that Theta -0

OpenStudy (anonymous):

\[\Theta \left(\begin{matrix} \\ 0\end{matrix}\right) =f(t) \neq constant\]

OpenStudy (anonymous):

yes true as theta=omega*time+1/2(acceleration*time^2)+theta_o where theta_o is at t=0:)